OpenStudy (anonymous):

HELP PLEASE Find the Solution Set then GRAPH 1.) 2x-3y=2 4x+5y=26 2.) x+y-6=0 3x-5y-34=0

OpenStudy (anonymous):

you can solve it by different methods, you can use matrices or simple substitution... using substitution method, 1)2x-3y=2 or x=(2+3y)/2 now put this expression for x in 4x+5y=26 and solve for y.. means solve 2(2+3y)+5y=26 for y

OpenStudy (anonymous):

then put value of y in either equation to find x

OpenStudy (anonymous):

i dont get it.

OpenStudy (anonymous):

i see...which grade are you in?

OpenStudy (anonymous):

tell me your grade.. so that i can explain accordingly

OpenStudy (anonymous):

8

OpenStudy (anonymous):

can you show me how to graph that

OpenStudy (anonymous):

you don't wanna know how to solve first?

OpenStudy (anonymous):

i wanna know

OpenStudy (anonymous):

ok lets do the first one 1.) 2x-3y=2 4x+5y=26